y"-4y'+13y=sin2t


1
Expert's answer
2021-06-15T09:22:52-0400

Solution.

y"-4y'+13y=sin2t

"\lambda^2-4\lambda+13=0"

"D=16-4\u202213=16-52=-36"

"\lambda=\frac{4\pm 6i}{2}=2\pm 3i"

From here solution of homogeneous equation "y''-4y'+13=0" is

"y=C_1e^{2t}\sin{3t}+C_2e^{2t}\cos{3t},"

where "C_1,C_2" are some constants.

Find particular solution of equation "y''-4y'+13=\sin{2t}" in the form "y=A\sin{2t}+B\cos{2t}."

Then

"y'=2A\cos{2t}-2B\sin{2t}," and

"y''=-4A\sin{2t}-4B\cos{2t}."


"(-4A+8B+13A)\sin{2t}+(-4B-8A+13B)\cos{2t}=\newline\sin{2t}"

Therefore, "A=\frac{9}{145},B=\frac{8}{145}."

So, we have solution

"y=C_1e^{2t}\sin{3t}+C_2e^{2t}\cos{3t}+\frac{9}{145}\sin{2t}+\frac{8}{145}\cos{2t}."

Answer. "y=C_1e^{2t}\sin{3t}+C_2e^{2t}\cos{3t}+\frac{9}{145}\sin{2t}+\frac{8}{145}\cos{2t}."